OpenStudy (anonymous):

When you add -10 to me and then subtract -8 the result is -16. What integer am I?

OpenStudy (ja1):

2? possibly

OpenStudy (anonymous):

Oh wait

OpenStudy (ja1):

34?

OpenStudy (anonymous):

x - 10 - 8 = -16 x = -16 + 10 + 8 x = 2

OpenStudy (anonymous):

It has to be a negative number

OpenStudy (anonymous):

is that in the question?

OpenStudy (anonymous):

yupp

OpenStudy (ja1):

-14 - 10 - (-8) = -16 -14 +( -10) + 8 = -16 -24 + 8 = -16 -16 = -16 so x = -14

OpenStudy (ja1):

this one was confuzzling :D

OpenStudy (ja1):

is -14 an answer?

OpenStudy (anonymous):

-14 is right

OpenStudy (ja1):

ok yay :D

OpenStudy (anonymous):

Thanks:)

OpenStudy (ja1):

no prob :)

OpenStudy (ja1):

I can haz medal?